OpenStudy (anonymous):

Solve square root of x-3 = square root of 2 subtract square root of x for...x Solutions: a.1/8 b.25/8 c.No solution d.-25/8

OpenStudy (tkhunny):

This? \[\sqrt{x-3} = \sqrt{2} - \sqrt{x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

start by squaring both sides. So, x-3 = 2 - 2sqrt(2x) + x Then simplify: 2sqrt(2x) = 5 -> sqrt(2x) = 5/2

OpenStudy (anonymous):

so is it b? which one is it?

OpenStudy (anonymous):

Can you take this last equation and square both sides?

OpenStudy (anonymous):

How? I don't know

OpenStudy (anonymous):

Yea I dont know

OpenStudy (anonymous):

sqrt(2x) = 5/2 -> 2x = 25 / 4, so x = 25/8

OpenStudy (anonymous):

Finally! Thank You

OpenStudy (anonymous):

You're welcome!